Pure acid is to be added to a 5% acid solution to obtain 95L of 45% solution. What amounts of each should she use?
Answer by Alan3354(69443)

Pure acid is to be added to a 5% acid solution to obtain 95L of 45% solution. What amounts of each should she use?
----------------------
100x + 5(95-x) = 45*95
95x + 475 = 4275
x = 40 liters of 100% acid